A juggler throws balls into air . He throws one whenever the previous one is at the heighest point how high do the balls rise if he throws ‘n’ balls each second? A projectile is fired from the ground with an velocity10m/s making an angle 60 with horizontal .how will be the path visibleto two observers, sitting in two buses moving with same speed along the same road in opposite direction,when speed of buses is [a] 10m/s [b] 5m/s [c] 2m/s

  1. A juggler throws balls into air . He throws one whenever the previous one is at the heighest point how high do the balls rise if he throws ‘n’ balls each second?
  2. A projectile is fired from the ground with an velocity10m/s making an angle 60 with horizontal .how will be the path visibleto two observers, sitting in two buses moving with same speed along the same road in opposite direction,when speed of buses is [a] 10m/s [b] 5m/s [c] 2m/s

Grade:11

1 Answers

Arun
5 years ago
Juggler throws the ball into the air with ball having  initial velocity = u
At the highest point of its path velocity will be zero as at maximum height velocity =0
So final velocity=0
According to question juggler throws a ball each second so for "n" balls 
time taken to reach highest position (t)= 1/n 
    v=u+gt
     0=u-g/n
    -u=-g/n
    u=g/n

S= ut-gt²/2
S=u/n-g/2n²
But u=g/n
S= g/n²-g/2n²
S=g/2n²